### v4

Abstract
DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome. This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics. Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences. Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences. Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully. We address this open problem by framing it as an “Embed-Search-Align” task. In this framework, a novel encoder model DNA-ESA generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment. In particular, DNA-ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. DNA-ESA is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem. DNA-ESA far exceeds the performance of $6$ recent DNA-Transformer model baselines such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
keywords:
Transformers, DNA Sequence Alignment, Large Language Models, Vector Stores
1 Introduction
Equal Contribution
Sequence alignment is a central problem in the analysis of sequence data. Many DNA sequencers generate short reads that are only a couple of hundred bases long. In order to interpret this data, a typical first step is to align the reads to a genome. Genomes come in many sizes, but many of the most commonly studied genomes, such as that of humans, are three gigabases long. Moreover, most experiments will generate millions of reads. Therefore, the resulting task of aligning reads to genomes is computationally expensive, and many decades of work have led to optimized approaches that can perform these tasks with great efficiency. However, many limitations remain with existing genome alignment methods. One limitation is that a reference genome usually is an approximation of the sequence from a single individual, and does not capture the variability of genomes across populations. Another, is that sequence alignment algorithms may not correctly model sequencing errors, and thus may introduce mistakes when trying to score the best match for a sequence against a genome. We asked whether some of these limitations could potentially be overcome if we utilize a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts. The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
We establish a foundation model foundation tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case). Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications. However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as genomes) and natural languages with large vocabulary. In LVLs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”. If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells. Existing foundation Transformer models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation. As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by. There are, however, tasks such as Sequence Alignment where the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance. Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective. This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LVLs to a continuous representation space.
The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads, where a sequencer generates a read of length $Q$
$r:=(\tilde{b}_{1},\tilde{b}_{2},\dots,\tilde{b}_{Q}),$ (1)
where $\tilde{b}_{i}\in\{A,T,G,C\}$. In practice, these reads come from individual genomes that do not necessarily match the reference and may contain mutations due to base insertions, deletions, and substitutions. Thus, it is assumed that this read is a noisy substring taken from a reference genome sequence $\mathcal{R}:=(b_{1},b_{2},\dots,b_{N}),\,b_{i}\in\{A,T,G,C\}$ and $N\gg Q$; for example, for the single-haploid human genome (chm_consort, ), $N\approx 3$ gigabases (gb), and $Q\approx 250$ though reads of much longer lengths are becoming increasingly affordable and accurate giab . The alignment task is to find a substring in $\mathcal{R}$
[…]
such that the edit distance $d(r,\tilde{r})$ – for example the Smith-Waterman (SW) distance – is minimized. The primary objective is to identify the most probable location, $q$, of this read within the reference genome.
Figure 1: Alignment Recall of Transformer-DNA Baselines by Read Length: Existing Transformer-DNA models were adapted for sequence alignment using mean-/max-pooling. Their performance, measured by recall (top-$K$) over $40K$ reads of varying lengths across the human genome (3gb - single-haploid), is shown. Trendlines represent each baseline, with error bars (Clopper-Pearson Interval (cps, ) @ 95%) in grey. The vertical line at $x=250$ marks a typical read length. Overall, these baselines show suboptimal performance. For more details, see Sec. 3.
Computational simulators have been developed to generate synthetic reads that have properties of real reads. These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (huang_art_2012, ).
((a)) Three $20,000$-long nucleotide sequences were selected, one each from Chr. 1, 2, and 3. Each sequence is broken up into $100$ consecutive reference fragments, each of length $1000$. Consecutive fragments have an overlap of $800$ bases (stride $200$). Additionally, reads of length $Q\sim\mathcal{U}[100,250]$ are sampled from within each fragment. The colors correspond to the respective chromosomes and the position of a particular fragment along the 20K-nucleotide sequence is coded by its color intensity.
((b)) Five gene regions (listed above in the inset) of different sizes are selected from the reference genome. Each gene is broken up into consecutive fragments of length $1000$ with a stride of $200$ (similar to subfigure (a)). Additionally, reads of length $Q\sim\mathcal{U}[100,250]$ are sampled from within each fragment. The colors correspond to genes and the position of a particular fragment along a specific gene is coded by its color intensity.
Figure 2: Illustrating DNA-ESA’s Preservation of Sequence Locality in Embedding Space: In DNA-ESA, the reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$. For effective sequence alignment, specific structures are expected in the embedding space: (1) Overlapping fragments $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ should have proximate embeddings; (2) Consecutive fragments forming a long sequence should correspond to a distinct manifold in the embedding space. Subfigures (a) and (b) display this emergent geometry, as visualized using DNA-ESA. In both subfigures, the fragment and read representations are jointly visualized in a 2D low-dimensional space constructed by applying UMAP umap to the original embeddings. Vertical marker lines (||||) refer to fragments, “–” markers correspond to reads. We observe that the consecutive fragments belonging to the same nucleotide (subfigure (a)) or gene (subfigure (b)) sequence constitute an order-preserving 1D manifold. Additionally, almost all of the reads are close to their corresponding fragments, making this space viable for the task of Sequence Alignment: by searching for fragments in the neighborhood of an external read represented in the embedding space, one is likely to retrieve a fragment most responsible for the read.
Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic SW approach, which computes the similarity between a read $r$ and reference $\mathcal{R}$; using the SW distance on its own, however, is computationally intensive for large genomes. Some key advances include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (bw_transform, ), reducing the effective search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (minimap, ; bwamem2, ; bowtie, ).
In this work, we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP). Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships. However, the introduction of models like the Transformers (vaswani_attention_2017, ) revolutionized this approach. Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ). The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020, ), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ). In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
1.1 Transformer models: Written Language to DNA Sequence Alignment
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences. Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, subsets of genes.
Within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function. Given these advances, a natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment? We delineate the associated challenges as follows:
  • [L1]
    Two-Stage Training: DNA-based Transformer models typically undergo pretraining via a Next Token/Masked Token Prediction framework, a method originally developed for natural language tasks. To form sequence-level representations, these models often employ pooling techniques that aggregate token-level features into a single feature vector. This approach, however, is sometimes critiqued for yielding suboptimal aggregate features  (reimers_sentence-bert_2019, ).
  • [L2]
    Computation Cost: The computational requirements for Transformer models grow quadratically with the length of the input sequence. This is particularly challenging for sequence alignment tasks that necessitate scanning entire genomic reference sequences.
Figure 1 shows the sequence alignment performance (recall) of several Transformer-DNA models. The testing protocols are elaborated in Sec. 3. Notably, these models exhibit subpar recall performance when aligning typical read lengths of 250.
2 Our Contributions
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder DNA-ESA, trained through self-supervision, to map DNA reads to relevant fragments in a reference sequence within a shared embedding space. (B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/. This provides efficient access to the entire reference sequence for each read alignment. These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
2.1 Task Definition
We formulate the problem of Sequence Alignment as minimizing a sequence alignment function, SA, applied to a read $r$ and a reference sequence $\mathcal{R}$ as
$v^{*}=\min_{q}\texttt{SA}(r,\mathcal{R})$ (3)
where $q\in\{1,2,3,\dots\}$ is a candidate reference starting position and $v^{*}$ is the optimal alignment score. Lower scores indicate better alignments. This optimization exhibits the following property:
Sharding for sequence alignment: For a read segment $r$ of length $Q$ and reference $\mathcal{R}$ of length $N$, the complexity of $\texttt{SA}(r,\mathcal{R})$ scales as $\mathcal{O}(NQ)$, which is expensive.
Ideally, we would like to use $\mathcal{R}$ to get a set of fragments $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$ that are a subset of the original reference. Then:
$v^{*}\approx\min_{\mathcal{F}_{i}\in\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}}\texttt{SA}\left(r,\mathcal{F}_{j}\right).$ (4)
Here, each $\mathcal{F}_{j}$ is a fragment of $\mathcal{R}$ (i.e., $\mathcal{F}_{j}\subset\mathcal{R}$), and $K$ is the number of these sub-tasks. This approximation is effective under the conditions:
  • (1)
    Fragment $\mathcal{F}_{j}$ lengths are on the order of the read length ($Q$), and not the length of the longer reference $\mathcal{R}$;
  • (2)
    There are enough fragments $\mathcal{F}_{i}$ to cover $\mathcal{R}$, i.e. $\cup\mathcal{F}_{j}=\mathcal{R}$;
  • (3)
    $K$ is significantly smaller than $\frac{N}{Q}$. If $\frac{N}{Q}$ then this amounts to scanning the whole reference.
Conditions (1) and (2) imply that fragments should be short and numerous enough to cover the reference genome. Condition (3) restricts the number of retrieved reference fragments per read — that we deem to be most likely to contain $r$ — to a small value $K$. Analogous methods have shown efficacy in text-based Search-and-Retrieval tasks  (peng2023, ; dai2022, ) on Open-Domain Question-Answering, Ranking among other tasks. Subsequent sections describe a parallel framework for retrieving reference fragments given a read. The pipeline is shown in Figs. 3 and 4.
Figure 3: System Overview [A] - Training Encoder and Populating Vector Store: Reference genome fragments $\mathcal{F}_{i}$ and within them, randomly sampled pure reads $r_{i}$ (positive pairs) are numerically represented via shared encoder $h$. Encoder training follows a contrastive approach as per (6). After training, the genome is segmented into overlapping fragments, encoded, and uploaded into the vector store.
2.2 Designing effective sequence representations
An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively – obey the following constraints over a pre-determined distance metric $d$:
$d\{h(r_{j}),h(\mathcal{F}_{i})\}\geq d\{h(r_{j}),h(\mathcal{F}_{j})\},\quad i\neq j$ (5)
Here $i$ and $j$ serve to distinguish whether a read is aligned to a particular reference fragment, a positive sample $\{r_{j},\mathcal{F}_{j}\}$, or there is a mismatch (negative sample): $\{r_{j},\mathcal{F}_{i}\}$. Observe that these inequalities constitute the only requirements for the encoder. As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed. Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment. This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
2.3 Self-supervision and contrastive loss
A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  chen_simple_2020 and gao_simcse_2021 : i.e. for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
$l_{r}=-\log\frac{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}}{e^{-d(h(r),h(\mathcal{F}_{j}))/\tau}+\sum^{B-1}_{i=1}e^{-d(h(r),h(\mathcal{F}_{i}))/\tau}}.$ (6)
Here $\tau$ is a tuneable temperature parameter. To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work simclr . Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
2.4 Encoder implementation
DNA-ESA uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks. The size of the vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. Generated numerical representations for each read, $h(r)\in\mathbb{R}^{1020}$. The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$. During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution. Here, $|x|$ denotes the length of $x$. To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch. Shorter sequences were padded to equal the length of the longest sequence in a batch. The distance metric used is Cosine Similarity.
2.5 Search and retrieval
Figure 4: System Overview [B] - Inference on a New Read: A read, as per (1) and generated by ART (huang_art_2012, ), is encoded by $h$. This is then compared to reference fragment representations in the vector database. The nearest-$K$ fragments in the embedding space are retrieved for each read, and the optimal alignment is determined using (7).
An outline of the search and retrieval process is presented in Fig. 4. Every read is encoded using the trained model and matched to reference fragments in the vector database. The top-$K$ retrieved fragments per read are then aligned using a SW alignment library to find the optimal alignment. The following sections describe the indexing and retrieval part in more detail.
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments :=$\mathcal{F}:=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots\}$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database. Once populated with all the fragments, we are ready to perform the alignment.
Retrieval: Given a read $r$, we project its corresponding DNA-ESA representation into the vector store and retrieve the approximate nearest-$K$ set of reference fragment vectors and the corresponding fragment metadata $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$.
Diversity priors: While the top-$K$ retrieved fragments can be drawn from across the entire vector store (genome), contemporary recommendation systems that use the top-$K$ retrieval setup rank and re-rank top search results (Slate Optimization – see grasshopper ) to ensure rich and diverse recommendations. Similarly, we apply a uniform prior wherein every retrieval step selects the top-$K$ per chromosome.
Fine-Alignment: A standard SW distance library (biopython, ) is used to solve (4), which can be executed concurrently across the $K$-reference fragments. Let the optimal fragment be $\mathcal{F}^{*}$. The metadata for each vector includes (a) the raw $\mathcal{F}^{*}$ sequence; (b) the start position of $\mathcal{F}^{*}$ within the reference $\mathcal{R}$, $q_{\mathcal{F}^{*}|\mathcal{R}}$. Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
$q^{*}=q_{|\mathcal{F}^{*}}+q_{\mathcal{F}^{*}|\mathcal{R}}.$ (7)
3 Transformer-DNA baselines
This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig. 1. We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(nucleotide, ), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(ji_dnabert_2021, ), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(nguyen_hyenadna_2023, ). Each model employs mean- and max-pooling of token representations for sequence encoding ($2\times 3=6$ baselines total). Independent vector stores for each baseline encode fragments from the entire $3$gb genome. We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}([25,1000])$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig. 1. Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective. These two baselines will be contrasted with DNA-ESA in Table 1.
DNA-ESA convergence plots are presented in Figs. 6(a) and  6(b). Model checkpoints are available at OSF. In Fig. 2, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (umap, ). The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in (6) encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance. (b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The DNA-ESA representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints. The codebase is linked.
4 Sequence alignment of ART-simulated reads
The results from Sec. 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see huang_art_2012 . ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
Simulator configurations: The different simulation configuration options and settings are listed: (A) Phred quality score $Q_{PH}$ in one of three $\{[10,30],[30,60],[60,90]\}$ ranges: the likelihood of errors in base-calls of a generated read; (B) Insertion rate $I\in\{0,10^{-2}\}$: the likelihood of adding a base to a random location in a read; (C) Deletion rate $D\in\{0,10^{-2}\}$: the likelihood of deleting a base in the read; (Others): Simulator system: MSv3 [MiSeq]; Read length: $250$.
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (4), and for this, we need to compute the SW distance. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$ – see (7) – as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
Distance bound, $d_{SW}\in\{1\%,2\%,5\%\}$: It is well known that short fragments frequently repeat in the genome and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled mappable ; reinventwillis . In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible. Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match. We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW distance between the pair ($v^{*}$ – see Eq. 4) and the optimal SW distance as a particular read length $Q$:
$d_{SW}=\frac{v^{*}-mQ}{(n-m)Q},$
where $m(=-2)$ is the match score and $n(=+1)$ is the mismatch penalty, hyperparameters in the computation of the SW distance. We consider an alignment (with $Q=250$) to be successful if the resulting SW distance between the read and the top returned fragment is within $d_{SW}$ of the optimal for that read length. A $d_{SW}=2\%$ (default) is equivalent to a mismatch of around $4$ bases in a read of length $250$.
Model Settings
ART
PacBio CCS
($\mu(v^{*}),\sigma(v^{*})$)
$I,D=0.01,Q=250$,
$d_{SW}=2\%$
Search [10,30] [30,60] [60,90]
Chr. 2, pbmm2
$d_{SW}=5\%$
Transformer Architectures HD (max) K=50 13.20 $\pm$ 1.81 17.01 $\pm$ 2.01 18.03 $\pm$ 2.05 -
DB2 (mean) 30.21 $\pm$ 2.44 39.19 $\pm$ 2.58 39.20 $\pm$ 2.58 -
DNA-ESA (ours) 98.40 $\pm$ 0.71 98.64 $\pm$ 0.67 98.60 $\pm$ 0.67 97.5 $\pm$ 0.82
K=75 98.80 $\pm$ 0.62 99.28 $\pm$ 0.52 98.88 $\pm$ 0.60
97.5 $\pm$ 0.82
(-498.0, 3.78)
Bowtie-2
- 99.80 $\pm$ 0.19
99.50$\pm$0.43
(-497.9, 3.78)
diff. - <1% <2%
Table 1: Performance of DNA-ESA with respect to baselines: The performance of DNA-ESA is compared to the DNA-BERT 2 and Hyena-DNA Transformer-based baseline models (with no additional finetuning) – the top performing baselines from Fig. 1. In addition, DNA-ESA is also compared with Bowtie-2 bowtie , a classical aligner. Comparisons are conducted across varying qualities (Phred score) of reads generated by ART huang_art_2012 . Additionally, an external PacBio CCS dataset of ($Q=250$ length-) reads from Chr. 2 of the Ashkenazim Trio - Son sample (as determined by pbmm2 minimap ) are aligned using both Bowtie-2 and DNA-ESA. All baseline models utilize a dedicated vector store. For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations. Across models and ART-generated datasets, the performance of DNA-ESA supersedes other Transformer-based models and is comparable to Bowtie-2 to within $1\%$. With respect to reads from PacBio CCS, DNA-ESA performs with $2\%$ of Bowtie-2 after controlling for the intrinsic quality of retrieved fragments according to the SW distance. Mean ($\mu$) and standard deviation ($\sigma$) across the top SW distance values ($v^{*}$) of reads successfully aligned in the case of DNA-ESA and Bowtie-2 models are reported.
4.1 Performance
Table 1 reports the performance of DNA-ESA across several read generation configurations – we choose from one of three ranges of Phred score $Q_{PH}$, $\{[10,30],[30,60],[60,90]\}$$I,D=0.01$, $d_{SW}=2\%$ – see Sec. 4), in addition to a direct comparison to DB2, mean and HD, max baselines (without any further finetuning), the best-performing baselines on pure reads identified in Sec. 3. In addition, we also compare DNA-ESA to Bowtie-2 bowtie , a conventional algorithmic aligner.
Additionally, models are also evaluated on an external set of reads from PacBio CCN generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) giab ). For (ii), the raw reads are $10$ kilobases long and for evaluation we consider random subset of $5000$ reads associated to Chr. 2333Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package minimap . Reads input into the aligner are cropped randomly to a length of $250$ to satisfy the computational requirements of DNA-ESA ($|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|=250$).
We observe that: (A) DNA-ESA shows strong recall performance of $>99\%$ across a variety of read generation and recall configurations; (B) On cleaner reads, DNA-ESA and Bowtie-2 are within $<1\%$ of each other in recall while controlling for the quality of the retrieved alignments (as determined by the SW distance). The results indicate that this constitutes a new state-of-the-art model for Transformer-based sequence alignment.
Sweeping Top-K and the SW distance bound In Table 2, we report the performance of DNA-ESA across distance bounds $d_{SW}\in\{1\%,2\%,5\%\}$ and a number of recalled fragment settings $K\in\{25,50,75\}$. Distance bound $d_{SW}<5\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments. Recall rates are comparable with different $d_{SW}$ values suggesting that when a read is successfully aligned, it is usually aligned to an objectively best match. Decreasing the top-$K$ per chromosome from $50\rightarrow 25$ does not substantially worsen performance ($<1\%$), indicating that the optimal retrievals are usually the closest in the embedding space. Additional ablation studies are presented in the Appendix.
$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
Top-25 / chromosome
0.0 0.0 97±0.94 98.12±0.76 98.56±0.69 97.24±0.91 97.96±0.0.80 98.6±0.69
0.01 96.88±0.97 98±0.78 98.36±0.73 97.24±0.91 97.96±0.80 98.76±0.64
0.01 0.0 97.08±0.93 98.04±0.78 98.56±0.69 97±0.94 98.24±0.75 98.92±0.59
0.01 97.16±0.80 97.96±0.56 99.2±0.52 97.6±0.85 98.12±0.76 98.76±0.62
Top-50 / chromosome
0.0 0.0 97.52±0.86 99.04±0.57 99.08±0.57 98.2±0.75 98.88±0.59 99.16±0.52
0.01 97.48±0.87 99.16±0.52 99.08±0.55 97.8±0.82 98.76±0.62 99.04±0.57
0.01 0.0 98.24±0.75 98.64±0.67 99.12±0.55 98.28±0.75 98.92±0.59 99.28±0.49
0.01 97.72±0.83 98.64±0.67 99.36±0.46 97.84±0.92 98.6±0.66 99.0±0.57
Top-75/ chromosome
0.0 0.0 98.04±0.78 99.12±0.55 99.0±0.57 98.4±0.71 99.12±0.55 98.4±0.71
0.01 98.4±0.71 99.04±0.56 99.4±0.46 98.44±0.71 98.84±0.62 99.24±0.52
0.01 0.0 98.64±0.67 99.0±0.57 99.2±0.52 98.0±0.78 98.72±0.64 99.48±0.43
0.01 98.68±0.64 99.28±0.49 99.4±0.46 98.48±0.69 98.88±0.59 99.4±0.46
Table 2: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec. 4. DNA-ESA presents a recall of $>99\%$ across several read configurations rivaling contemporary algorithmic models such as Bowtie. As expected, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
4.2 Task transfer from Chromosome 2
Are these results indicative of DNA-ESA’s adaptability to new genomic sequences rather than strict adherence to its training data? This would suggest the model’s learning to solve the sequence alignment task rather than memorizing the genome.
[…]
Acc. $\uparrow$ F1 $\uparrow$ P $\uparrow$ R $\uparrow$
Train ART Reads 2 (seen in training) 3 (populated) Y (populated)
Chr. 2 $Q_{PH}\in[30,60]$ 99.74 $\pm$ 0.29 99.74 $\pm$ 0.29 $9$$\pm$ 0.01 99.49±0.46 98.87±0.59 99.99±0.01
Chr. 2 $Q_{PH}\in[60,90]$ 99.37 $\pm$ 0.49 99.36 $\pm$ 0.49 $9$$\pm$ 0.01 98.74±0.64 99.12±0.55 99.5±0.43
Table 3: DNA-ESA Performance on Chr. 2 and Task Transfer to Chr. 3, Y: DNA-ESA, trained on Chr. 2, is evaluated for aligning positive samples from Chr. 2 and negative samples from Chr. 3. Anticipated failure to align Chr. 3 reads (false positives) is due to significant SW distance. Metrics reported include weighted accuracy (Acc.), precision (P), F1-score (F1), and recall (R), for comparison with Table 1. [Grey] For Chr. 3 and Y, separate vector stores are populated using DNA-ESA representations learned using Chr. 2 {read,fragment} pairs. These results support the generalizing tendency of DNA-ESA to align sequences beyond the training reference.
Species $Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
Recall Top-K @50 $\uparrow$
Thermus Aquaticus [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
Acidobacteriota [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
Rattus Norwegicus [Chr. 1] 95.98±1.07 96.48±1.01
Rattus Norwegicus [Chr. 2] 97.99±0.78 96.49±1.01
Pan Troglodytes [Chr. 2A] 98.5±0.69 95.74±1.11
Pan Troglodytes [Chr. 2B] 95.99±1.08 96.25±1.05
Table 4: Cross-Species Task Transfer with DNA-ESA: Training on Human Chr. 2, Testing on Diverse Species: DNA-ESA, trained on human Chr. 2, aligns fragments and reads from different species, including Rattus Norwegicus, Pan Troglodytes, Acidabacteriota, and Thermus Aquaticus. These species, are evaluated for read alignment recall using ART-generated reads. The findings, consistent with Table 1, indicate DNA-ESA’s proficiency in modeling DNA sequence structure, beyond simply memorizing training data.
Performance: Details on convergence are in Appendix 8. Table 3 shows that the performance on unseen human chromosomes (3, Y) is similar to the performance reported in Table 1, despite only training DNA-ESA on Chr. 2. This suggests DNA-ESA’s ability to generalize sequence alignment across chromosomes with different compositions. High Accuracy, Precision, and F1-scores further confirm the method’s specificity at this high recall. In Table 4, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (originstory, ). Even distantly related species like Thermus aquaticus and Acidobacteriota show significant recall, highlighting DNA-ESA’s task transferability beyond simple data memorization.
5 Limitations and Future Work
DNA-ESA has the following challenges that we hope to overcome in future generations of the system: the computational speed, at 10K reads per minute, is a limiting factor for extensive genomic studies involving millions of reads. Despite this, the alignment task’s inherent parallelizability offers avenues for efficiency improvements. Our ongoing and future efforts are directed towards overcoming this time efficiency bottleneck, as detailed in Appendix Sec. 9. We are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. Additionally, we aim to enhance DNA-ESA’s performance at shorter read lengths, exploring alternative training methods and integrating more diverse read/reference features.
6 Concluding Remarks
Current DNA sequence alignment methods, honed over decades, incorporate DNA-specific enhancements in both indexing and retrieval, relying on algorithmic approaches. However, we introduce an alternative data-driven paradigm, employing a Transformer-based DNA-ESA encoder. Our method performs sequence alignment through self-supervised learning, utilizing contrastive loss. While similar methods have previously found application in identifying approximate semantic similarity in written language, our approach surprisingly excels at identifying exact overlaps among DNA sequences. Empirical results underscore the model’s ability to embody the inherent structure of any DNA sequence, irrespective of location or species origin. A distinguishing feature of DNA-ESA is its ability to map both shorter reads and longer reference fragments into the same space by employing cosine similarity as a symmetric distance metric. This complements existing paradigms that have traditionally used a hierarchical and explicit comparison of base distributions.
DNA-ESA representations would find utility in other related tasks. If, for example, one had to align reads to the Pan Genome liao instead of a single reference, one can populate a single vector store with fragments from all $47$ genome assemblies. Then given a read, the nearest fragment to the read can belong to any reference variant. Note that in this case, we avoid having to construct a separate index for each of the references as is usually done in conventional methods.
One can similarly repurpose the representation space – originally intended for Sequence Alignment – to address the complementary task of Genome Assembly. We conclude this paper with a setup that provides initial proofpoints under reasonable simplifying assumptions about read lengths and overlaps among the reads. The resulting framework will be formalized and benchmarked with reads from actual experiments in our future work.
6.1 A setup for de-novo Genome Assembly using DNA-ESA
Consider a set of reads, $R:=\{r_{1},r_{2},\dots,r_{C}\}$ from which we would like to estimate a reference $\mathcal{R}$. The de-novo assembly task is to estimate an ordering of the reads in $R$, $\hat{R}:=(r_{i},\dots,r_{j},\dots,r_{k})$ such that a concatenation (after removing overlaps) is approximately the original reference. In typical setups, all pairs of reads would need to be compared in order to find those that overlap and construct longer chains of the assembly in an incremental fashion. This may be visualized as a complete graph $K_{C}$ with $C$ nodes (reads) and $C^{2}$ edges (constraints). To avoid the quadratic complexity, modern techniques such as Overlap-Layout-Consensus (OLC) or de Bruijn Graph (DBG) approaches assembly have adopted hot-start heuristics such as base overlap and shared K-mer counts to reduce the number of pairwise read comparisons.
The representation space modeled by DNA-ESA admits a comparable assembly setup. Fig. 2 suggested emergent 1D manifolds in the representation space – comprising the embeddings of the reads – $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ – such that the relative positions of fragments along the manifold correlated to the relative locations of the fragments along the reference. Therefore, a pairwise read-read comparison within a smaller radius in the embedding space would be sufficient to find reads that overlap. And more generally, a walk along the 1D manifold would constitute a near-optimal assembly.
Figure 5: Motivating the task of de-novo assembly using the distance adjacency matrix of the UMAP-generated $k$-nearest neighbor ($k=10$) network of reads from Thermus Aquaticus: We assume that reads $r_{1},r_{2},\dots,r_{C}$ belonging to a species are embedded into the DNA-ESA embedding space $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ using a model trained on a known reference genome belonging to a different species. For this illustrative numerical example we ensure that reads are created with the following properties: first, no read is a substring of any other read (efficient algorithms for the case where reads can be substrings of others will be covered in our future work); second every read has a corresponding read that it overlaps with for each end, and finally, the union of the reads cover the ground-truth reference genome. Thus, for every read $r_{i}$ there exist reads $r_{k}$ and $r_{j}$ such that the ordered triplet $(r_{j},r_{i},r_{k})$ constitutes a fragment of the reference. In a perfect embedding scenario, the closest neighbors of $r_{i}$ (i.e. $k=1$ in an $k$-NN search) in the embedding space should be $r_{j}$ ad $r_{k}$. For this ideal case, if a network is constructed where each read (node) is connected to its two nearest neighbors, and the nodes are arranged based on the order in which they appear in the reference chromosome, one would get a bi-diagonal adjacency matrix. When ground truth is not known, the adjacency matrix will be a permuted version of a bi-diagonal matrix, but a complete assembly can still be performed by a greedy algorithm that starts with any node and pieces together the neighboring reads in the embedded space in time that grows linearly in the number of reads. Simulated reads, each of length $500$ and with random overlaps of lengths drawn uniformly $\mathcal{U}([150,250])$ with its neighboring reads at each end, are sampled from $4$ of the chromosomes in Thermus Aquaticus, and are embedded into the DNA-ESA embedding space trained with a reference human genome. The sorted adjacency matrix of the UMAP umap -generated $k$-NN ($k=10$) network is presented above for each chromosome. The dark off-diagonal distance values in the $k$-NN adjacency graph with $k=10$ show that the distance properties of actual embeddings of simulated reads from Thermus Aquaticus are almost coincident with that of the ideal case discussed above. Moreover, just as in the ideal case, by finding an approximate solution to the Traveling Salesman Problem in the unsorted network, one can find a walk (the assembly) that is close to the original reference. The results are presented in Table 5.
We provide an initial computational implementation of this idea: (a) The k-nearest neighbor (kNN) graph $G(R,d)$, representing the distances between close-by reads, is computed using the density-aware corrections provided by the UMAP library umap : graph adjacency matrices for each of the chromosomes is visualized in Fig. 5. For this figure, the several reads are ordered manually as they appear in the assembly in order to highlight the connectivity pattern; the bright off-diagonal values indicate a chain-like structure corresponding to the observed manifold; (b) A Hamiltonian cycle computed through the graph that has the lowest total distance corresponds to an assembly, $\hat{R}$. This involves solving the Traveling Salesman Problem (TSP) – we use the 3/2-approximate Christofides algorithm christofides . After a walk across the reads is generated, a second pass through the walk greedily concatenates adjacent reads $(r_{i},r_{i+1})$ by computing the alignment using the SW distance and removing duplicates and overlaps. In cases where adjacent reads do not have any alignment to one another – the walk isn’t guaranteed to be perfect since the DNA-ESA model is heuristical and the TSP algorithm is approximate – the current assembly part is stowed away as a contig contig , and a new contig is instantiated with the incoming read. The fewer the contigs, the better the manifold in the representation space correlates to the ordering of the true reference.
Assembly is conducted across $4$ of the shorter chromosomes in the Thermus Aquaticus genome. We consider (long) reads of length $Q=500$, which have no indels, and the entire corpus of reads is ensured to have full coverage across each chromosome. Each read has an overlap of $\mathcal{U}([150,250])$ bases with another read in the corpus of reads. QUAST quast metrics are reported in Table 5.
Size (in kb) $\uparrow$ No. of Contigs $\downarrow$ N50 (in kb) $\uparrow$ NGA50 (in kb) $\uparrow$
Thermus Aquaticus NZ_CP010823.1 14.4 1 14.4 14.4
NZ_CP010824.1 16.6 1 16.6 16.6
NZ_CP010825.1 69.9 22 6.0 4.4
NZ_CP010826.1 78.7 15 10.3 10.3
Table 5: DNA-ESA performance on de-novo assembly of the Thermus Aquaticus genome: The solution to the approximate TSP applied on the DNA-ESA read-read network generates a set of contigs that can be evaluated with respect to the original reference using QUAST quast . For the two shorter chrosomomes, the resulting assembly was perfect; the entire reference is encapsulated by one contig. Even in the case with longer chromosomes, the number of contigs are still few and of significant length (as indicated by the high N(G)50 score) suggesting a strong correlation between the distance metric in the embedding space and the ordering of the corresponding reads in the assembly.
N50 refers to the length of a contig (in bases) such that the contigs that exceed this length span 50% of the reference genome. NGA50 is a reference-aware N50 metric. The fewer contigs, large N(GA)50 scores in comparison to the large size of the chromosomes, when considered together, demonstrate that the reads are indeed ordered in a relative position-aware manner in the DNA-ESA embedding space with respect to a reference making the de-novo assembly task viable.
{appendices}
7 Appendix
8 A. Convergence
Fig. 6 plot the convergence of the DNA-ESA encoder model discussed in the main text. We see convergence after $\sim$20k steps.
((a)) Trained on the Human Genome
((b)) Trained on the Human Chromosome 2
Figure 6: DNA-ESA convergence plots. Both plots show the loss pr. step (grey). For clarity, we smooth the loss using a moving average (black).
9 B. Complexity of computing alignment
9.1 Cost of constructing a new representation
Computing the embedding $\mathcal{E}$ of a sequence of length $F$ using DNA-ESA encoding – a typical Transformer-based attention architecture – has the following computation complexity:
$\mathcal{O}(LH*(F^{2}*d+d^{2}*F))\Rightarrow\mathcal{O}(F^{2}*d+d^{2}*F)$ (8)
Where $d$ is the embedding dimension of the model, $L$ is the number of layers in the Transformer and $H$ is the number of heads per layer. As $d$ is a controllable parameter for the model, we can further simplify:
$\mathcal{O}(F^{2}*d+d^{2}*F)\Rightarrow\mathcal{O}(F^{2})$ (9)
The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (beltagy_longformer_2020, ; kitaev_reformer_2020, ) have developed shortcuts to reduce the cost. These have already been applied to DNA sequence modeling  (nguyen_hyenadna_2023, ).
9.2 Vector Store Upstream
Vector store $\mathcal{D}$ is populated once (in bulk) with encoded fragment-length sequences drawn from the entire genome; constant time complexity $C$ to upload $<10M$ vectors.
9.3 Retrieval Cost
Given a new embedding, $\epsilon(G,K)$ is the cost of retrieving top-$K$ nearest neighbors across the fragment embeddings, where $G$ is the length of the reference genome. In modern vector databases, where several hashing techniques such as approximate K-nearest neighbors are used, $\epsilon$ scales logarithmically with $G$. This is indeed the key benefit of using such databases.
9.4 Fine-grained Alignment
Existing libraries/algorithms (e.g. the Smith–Waterman algorithm) can identify the alignment between a fragment sequence (of length $F$) and read (of length $Q$) in $\mathcal{O}(FQ)$.
9.5 Total Complexity
Total complexity involves (a) constructing the representation of a read; (b) querying the vector store; (c) running fine-grained alignment with respect to the $K$ returned reference fragment sequences:
$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))$
$\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
10 C. Ablation studies
10.1 Without diversity priors in top-K
In Table 6, we report the performance without diversity priors used in the retrieval step: i.e. the nearest-K neighbors in the embedding space are selected from the entire set of fragments spanning the genome rather than uniformly sampling from each chromosome. The performance predicably falls in comparison to those reported in Table 1 since fewer fragments scattered unevenly across the different chromosomes are being retrieved per read.
$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
[…]
0.0 0.0 98.56 $\pm$ 0.67 97.60 $\pm$ 0.85 98.56 $\pm$ 0.67 97.76 $\pm$ 0.82 97.76 $\pm$ 0.82 97.44 $\pm$ 0.88
0.01 0.01 98.8 $\pm$ 0.62 98.4 $\pm$ 0.71 98.8 $\pm$ 0.62 99.2 $\pm$ 0.52 98.4 $\pm$ 0.71 98.4 $\pm$ 0.71
Top-50 global
0.0 0.0 91.6 $\pm$ 1.49 92.4 $\pm$ 1.43 93.8 $\pm$ 1.31 90.4 $\pm$ 1.58 93.8 $\pm$ 1.31 94.8 $\pm$ 1.21
0.01 0.01 89.6 $\pm$ 1.64 91.6 $\pm$ 1.49 96 $\pm$ 1.07 94.4 $\pm$ 1.25 94.4 $\pm$ 1.25 93.2 $\pm$ 1.36
Table 6: Sequence alignment recall of DNA-ESA - without diversity priors: The various parameters are described in Sec. 4. When compared to the main results presented in Table 1, it is evident that the addition of diversity priors results in an improvement of $\sim$ in recall. Similar to the result presented in the main text, performance improves with larger search radius in the vector store (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
10.2 Across read lengths
In Table 7, the performance of DNA-ESA is reported across read lengths. We observe Zero-shot performance at longer read lengths: The model performs better at longer read lengths (even exceeding the read length bound established during training $\mathcal{U}[150,500]$); while evaluating for longer reads, we make sure to guarantee that the reads exist as subsequences of fragments. Improving the performance at shorter read lengths is the subject of future work.
I D $Q=200$ $Q=225$ $Q=250$
$Q_{PH}\in[30,60]$
0 0 99 $\pm$ 0.57 98.6 $\pm$ 0.67 99 $\pm$ 0.57
0.01 0.01 97.4 $\pm$ 0.88 98.8 $\pm$ 0.62 98.6 $\pm$ 0.67
$Q_{PH}\in[60,90]$
0 0 98.6 $\pm$ 0.67 98.6 $\pm$ 0.67 98.9 $\pm$ 0.59
0.01 0.01 98.6 $\pm$ 0.67 99 $\pm$ 0.57 98.6 $\pm$ 0.67
Table 7: DNA-ESA recall performance across read lengths: Performance of DNA-ESA is higher for longer reads including those lengths on which the model was not trained ($Q\in\mathcal{U}([150,500])$). Shorter reads are more challenging for the model potentially due to replicates found across the reference.
11 Competing interests
No competing interest is declared.
12 Author contributions statement
VR devised the task and framework and assisted with writing. MP aided in writing and bridging NLP and Bioinformatics communities. LSB contributed to writing and brainstorming. PH and KCE realized the framework and assisted with writing. SR ran several experiments to improve model performance. LM and TG reviewed existing conventional baselines.